# MaskINT: Video Editing via Interpolative Non-autoregressive Masked Transformers

### v1

@@ 7 Conclusion
We propose MaskINT towards consistent and efficient video editing with text prompt. MaskINT disentangle this task into keyframes joint editing with diffusion methods and structure-aware frame interpolation with non-autoregressive masked transformers. Experimental results demonstrate that MaskINT achieves comparable performance with pure diffusion-based methods while significantly reduces the inference time. Moreover, our work demonstrates the substantial promise of non-autoregressive generative transformers within the realm of video editing.
@@ 8 Diverse Structural Conditions
Table 4: Quantitative comparisons of the combination of varied structural conditions in each stage on ShutterStock. “T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.

### v2

@@ 7 Conclusion
We propose MaskINT towards consistent and efficient video editing with prompt. MaskINT disentangles this task into keyframes joint editing with T2I model and structure-aware frame interpolation with non-autoregressive masked transformers. Experimental results demonstrate that MaskINT achieves comparable performance with pure diffusion-based methods while significantly accelerate the inference up to 7x. We believe our work demonstrates the substantial promise of non-autoregressive generative transformers within the realm of video editing.
@@
1 Dataset Details
ShutterStock [1] is a commercial dataset with over 30 million paired text-video examples. Our model is trained using a subset of 100k videos. The WebVid dataset [2] offers an accessible version of ShutterStock, but they include a central watermark due to copyright constraints.
@@ 2 Diverse Structural Conditions
Table 1: Quantitative comparisons of the combination of varied structural conditions in each stage on ShutterStock. “T.C." stands for “temporal consistency", and “P.C." stands for “prompt consistency".